A COUPLED CLUSTER STUDY OF LINEAR C_{7}: NO EVIDENCE OF FLOPPINESS

\underline{PETER BOTSCHWINA}, Institut für Physikalische Chemie, Universität Göttingen, Tammanstraße 6, D-37077 Göttingen, Germany.

Linear C_{7} has been studied by means of coupled cluster calculations including connected triple substitutions (variant CCSD(T)) and a large basis set of 385 contracted Gaussian type orbitals. The calculations indicate that linear C_{7} is a fairly normal semi-rigid molecule with no sign of floppiness. In particular, excitation of the bending vibration with lowest wavenumber (ν_{11}) produces only a 0.2% change in the rotational constant. The results of the present calculations are at variance with an experimental study by Heath and Saykally^{a}. A renewed analysis of the experimental data and possible remeasurement is strongly recommended. Rovibrational transitions within the ν_{11} band of linear C_{7} are predicted around 70 cm^{−1} and thus rather close to transitions within the ν_{2} band of non-rigid C_{3}, but with very different spacing.

aJ. R. Heath and R. J. Saykally , J. Chem. Phys. \underline{94}(1724), 1991.